# A Scalable Unsupervised Framework for multi-aspect labeling of Multilingual and Multi-Domain Review Data

### v1

@@ 1 Introduction
User review data contains direct feedback from consumers about products and services, and can be used as an important source of information for business decision-making, marketing strategy development, and brand reputation management. However, most review data exists in unstructured text format, and typical sentiment polarity analysis or star rating–based analysis has the limitation of not sufficiently capturing the detailed attributes embedded in the reviews. As an alternative to overcome these limitations, Aspect-based Analysis has been gaining attention. Aspect-based Analysis can be widely used in various fields such as consumer feedback analysis, public policy monitoring, and social media analysis (Khabiri et al., 2021; Qiu et al., 2011), and has the advantage of enabling detailed understanding of how each aspect is evaluated in text.
However, even Aspect-based Analysis has several limitations when applied to user review data, which can be broadly categorized into four aspects. First, most studies on Aspect-based Analysis focus on specific domains such as electronics or restaurant reviews (Zhao et al., 2020), making it difficult to ensure that the proposed methods show the same performance and validity in real service environments where a wider range of domains must be addressed. In other words, studies conducted within limited domains do not sufficiently demonstrate multi-domain applicability, which restricts their practical use across industries. Second, Aspect-based Analysis research tends to focus on quantitative metrics such as accuracy and F1-score (Zhang et al., 2024), and lacks qualitative research verifying whether the aspects extracted by the model are semantically valid and practically effective in the real world. As a result, it becomes difficult to determine whether the extracted aspects are trustworthy enough to be used for problem diagnosis or product improvement in practice, which again acts as a limitation in applying the methods in industry. Third, most Aspect-based Analysis research adopts approaches that fine-tune labeled data or assume specific data formats, which leads to high labeling costs and necessitates format conversion (Mukherjee et al., 2021). When relying solely on labeled data, it becomes difficult to accurately capture aspects from newly appearing review texts that are not labeled, leading to limitations in responding in real time in actual service environments. Lastly, another issue is that most Aspect-based Analysis studies are conducted mainly on English data. Since large-scale review data is accumulated in various languages in global service environments, models trained only in a specific language face limitations in expanding to multilingual data. As a result, reviews written in newly emerging or underrepresented languages may not be sufficiently analyzed, leading to missed opportunities for satisfaction analysis and service improvement in various markets.
@@ 2.1 Aspect-based Analysis
Research focusing on the analysis of specific aspects in various types of text data has been actively conducted. This technique is utilized in diverse fields such as consumer review analysis, public opinion monitoring, stock market trend prediction (Lee et al., 2018b, a; Park and Kim, 2019), and online marketing strategy development (Mukherjee et al., 2020; Angelidis et al., 2021).
@@
Early studies mainly relied on manually defined rule-based approaches, applying topic modeling methods such as Latent Dirichlet Allocation (LDA) to extract aspect terms from text (Khabiri et al., 2021). Qiu et al. (2011) proposed extracting aspects using part-of-speech tagging and word frequency data. Another method involved setting preprocessing rules, performing frequency analysis on text data, and manually assigning keywords to predefined categories based on domain knowledge to evaluate keyword importance per attribute (Shofwatalloh et al., 2024). Zhao et al. (2020) proposed a method to extract aspect terms and opinion terms simultaneously, but since it mainly focused on identifying individual elements, it had limitations in effectively extracting aspect-opinion pairs. These approaches depend on predefined rules, making them difficult to apply to various domains, and because aspect categories are selected based on domain knowledge, they have limitations in reflecting domain-specific terms. Moreover, since most analyses are limited to nouns, it is difficult to conduct sentiment analysis that reflects context.
@@
To overcome these limitations, supervised learning–based approaches have been introduced. Early studies explored the use of Conditional Random Fields (CRFs) to simultaneously extract structural features of sentences and polarity opinions (Li et al., 2010). More recently, deep neural network–based approaches have shown promising performance. Wang et al. (2016) applied an attention-based Long Short-Term Memory (LSTM) model for aspect-level sentiment classification, allowing the model to focus on important words in the sentence based on the input aspect. Fan et al. (2019) defined a method called Target-oriented Opinion Words Extraction (TOWE) using Inward-Outward LSTM to extract opinion words related to specific targets. Additionally, the RNCRF model, which combines Recurrent Neural Networks (RNNs) and CRFs, was introduced to learn the relationship between aspects and opinions, and a joint learning method was proposed by optimizing CRF parameters (Wang et al., 2016). Topic-level sentiment analysis research has also been conducted. Approaches using Online Latent Semantic Indexing and LSTM-based models have been proposed to classify sentiment for specific topics in sentences (Pathak et al., 2021). Mikolov et al. (2013) conducted research on learning word representations using neural network–based language models. However, supervised deep learning models require a large amount of labeled data, which involves high cost and time for data construction. Moreover, models trained in a specific domain are difficult to apply directly to other applications.
@@
To address these limitations, recent studies have drawn attention to unsupervised learning–based approaches, which aim to perform effective aspect extraction and sentiment analysis without labeled data. Local topic models at the sentence level have been proposed to automatically infer aspects and generate seed sets of sentiment words without manual annotation (Brody and Elhadad, 2010). Mukherjee and Liu (2012) proposed a semi-supervised topic model that jointly models aspects and sentiment. Zhao et al. (2010) combined topic modeling and maximum entropy (MaxEnt) models to separate aspect and opinion words using a semi-supervised approach. More recently, unsupervised aspect term extraction using attention-based Deep Neural Networks (DNNs) has been proposed (He et al., 2017). These approaches have the advantage of automatically clustering semantically similar aspects without labeled data. However, existing unsupervised methods lack sufficient verification of cross-domain generalization performance, and due to instability during training, they often suffer from low reproducibility and inconsistent performance.
@@
Most previous studies have been conducted in single-language settings. For example, in Aspect-Based Sentiment Analysis using Korean data, Park and Shin (2020) explored Attribute Category Sentiment Classification (ACSC) using Bidirectional Encoder Representations from Transformers (BERT). They replaced the [CLS] token with output vectors from tokens related to attribute categories and analyzed performance differences according to input configurations (QA/NLI). In a study using beauty commerce reviews, Importance-Performance Analysis (IPA) and DistilBERT-based sentiment analysis were used to identify areas for product improvement (Shofwatalloh et al., 2024). Another study analyzed sentiment for specific evaluation categories such as food, price, service, and atmosphere using Korean restaurant reviews and proposed a model for predicting detailed ratings (So and Shin, 2020). Lee and Ahn (2022) applied attribute-based sentiment analysis using BERT in a movie recommendation system to classify reviews by attributes (e.g., direction, actors, story) and develop a Movie Multi-Criteria recommendation model reflecting user preferences. These studies demonstrate that Aspect-Based Sentiment Analysis enables more fine-grained evaluation of products and services and can enhance the performance of recommendation systems. However, they still lack sufficient verification of generalization across domains and studies considering multilingual applicability remain scarce.
@@ 2.2 Attention Mechanism in Aspect-based Analysis
Recently, attention mechanisms have been widely used in Natural Language Processing (NLP), significantly enhancing the performance of deep neural network models. In particular, pre-trained models such as BERT, Robustly Optimized BERT Pretraining Approach (RoBERTa) effectively capture contextual information through attention, achieving state-of-the-art performance in various NLP tasks. Attention mechanisms learn by focusing on important information within the text. For example, in sequence tagging tasks, they highlight specific words within context (Li et al., 2021), and in sequence-to-sequence learning, they help maintain core information during sentence transformation (Liu et al., 2019). In sentiment analysis, attention helps models perform more precise predictions by focusing on words related to specific aspects (Wang et al., 2021), and in machine translation, it contributes to improving semantic alignment between source and target languages (Zhang et al., 2023). Many researchers have explored the use of attention mechanisms in supervised models for effective Aspect-Based Sentiment Analysis (ABSA) (Xu et al., 2019). Additionally, unsupervised attention-based models have been proposed to extract aspects from text without large-scale labeled data (He et al., 2017). However, most of these existing studies focus on single languages and specific domains, and there is still a lack of research verifying generalization performance in multilingual and multi-domain environments.
@@ 3 Dataset
For Korean review data, we used the Korean String Processing Suite (KSS) [^1] https://github.com/hyunwoongko/kss to segment the text into individual sentences for sentence-level analysis. As a result, we obtained 313,333 sentence-level entries for hotel reviews and 288,874 for food reviews. After removing unnecessary special characters and extra spaces, we refined the text using the Open Korean Text (OKT) morphological analyzer from Konlpy (Park and Cho, 2014). Since Korean features various grammatical transformations including verb and adjective inflections, we lemmatized these parts of speech and removed case particles (e.g., Korean case particles such as i/ga (subject), eul/reul (object), e (to/at), eseo (from/in)) to ensure analytical consistency. For example, the Korean expressions “맛있었다” (delicious + past tense) and “맛있을 것이다” (delicious + future tense) were both normalized to “맛있다” (delicious). This preprocessing step was critical for reducing variability caused by Korean-specific tense and ending changes that could interfere with model training. We also customized the OKT user dictionary by adding neologisms, slang, and domain-specific proper nouns. This helped prevent incorrect token splits of words not included in the default dictionary and improved the performance of the analyzer. Both preprocessed and original sentences were retained to enable flexible follow-up analyses without data loss.
@@ 4.1.1 Embedding Vector Extraction
We train word embeddings reflecting semantic similarity using the Word2Vec model (Mikolov et al., 2013). Specifically, we apply the Continuous Bag of Words (CBOW) approach, which predicts a center word based on its surrounding context words. For example, in the sentence “The hotel has great service and staff”, if “service” is the center word and the window size is 1, the context words are “great” and “and”.
@@ 4.2.1 Multi-Head Attention-Based Sentence Representation
Attention mechanisms learn the relationships between words in a sentence and reflect the importance of each word in context. Unlike earlier studies (He et al., 2017) that relied only on attention, we use Multi-Head Attention to learn sentence representations that capture contextual information.
@@
$W^{O}$ is the output weight matrix that projects the concatenated outputs of all attention heads into the final output dimension. Here, $head$ denotes an individual attention head, and each attention head is computed as shown in Eq. 4.
@@ 4.2.2 Aspect-Aware Embedding Vector
In this study, the aspect-aware embedding vector $r_{i}$ is computed as shown in Eq. 5, by leveraging the aspect embedding matrix $E_{c}$ calculated in Section 4.1 and the probability vector $p_{t}$, which represents the likelihood of the sentence being associated with each aspect category.
@@ 4.2.3 Optimization of Aspect-Aware Representations
In this study, we adopt the Max-Margin Loss to train the aspect-aware embedding vector $r_{i}$ introduced in Section 4.2.2. In neural language models, the negative sampling technique is widely used to encourage the model to increase the similarity between a sentence and its corresponding aspect category, while reducing the similarity with negative samples. During the application of the negative sampling method, the model is guided to learn more generalizable patterns by minimizing the inner product between the sentence vector $z_{i}$, generated in Section 4.2.1, and negative samples, while maximizing the similarity with the most relevant aspect. The negative samples $z_{n}$ are composed of representation vectors of sentences unrelated to the input sentence, which helps prevent overfitting to a specific aspect during training. To achieve this, we define the hinge loss function as shown in Eq. 6.
@@
[…]
@@ 5 Experimental Result
This section comprehensively evaluates the performance and practicality of the proposed MUSCAD framework through a series of experiments. First, we empirically determined the optimal number of clusters for each domain during the construction of the automatically labeled dataset using the framework. The number of aspect categories per domain was also determined through expert review. Then, we conducted multi-label classification experiments using the constructed datasets to compare the performance of fine-tuned classification models with that of LLM–based labeling. In addition, a Human Evaluation was conducted to assess whether the automatically generated labels are semantically valid in real-world domain settings. Through this, we verify the reliability and scalability of the proposed framework.
@@ 5.1 Experimental Settings
To validate the quality of the labeled datasets generated by the framework, we fine-tuned several pre-trained language models and evaluated their classification performance. We randomly undersampled 40,000 samples for each domain. For the Korean datasets (Hotel and Food), we used KoELECTRA (Park, 2020), KR-BERT (Lee et al., 2020), KLUE-BERT (Park et al., 2021), Multilingual BERT (Pires et al., 2019), and XLM-RoBERTa (Goyal et al., 2021). For the English dataset (Beauty), we used Multilingual BERT, RoBERTa (Liu et al., 2020), BERT-base, BERT-large (Devlin et al., 2019), XLM-RoBERTa, and DeBERTa-base (He et al., 2021). All pre-trained language models were fine-tuned using the same hyperparameter settings: a batch size of 16, a maximum token length of 128, and a learning rate of 5e-5. We used the AdamW optimizer and trained each model for 5 epochs. During training, we applied 5-fold cross-validation to ensure stable performance across the dataset and evaluated the models using both Micro-F1 and Macro-F1 scores.

### v2

@@ 1 Introduction
User review data contains direct feedback from consumers about products and services, and can be used as an important source of information for business decision-making, marketing strategy development, and brand reputation management. However, most review data exists in unstructured text format, and typical sentiment polarity analysis or star rating–based analysis has the limitation of not sufficiently capturing the detailed attributes embedded in the reviews. As an alternative to overcome these limitations, Aspect-based Analysis has been gaining attention. Aspect-based Analysis can be widely used in various fields such as consumer feedback analysis, public policy monitoring, and social media analysis [6, 7], and has the advantage of enabling detailed understanding of how each aspect is evaluated in text.
However, even Aspect-based Analysis has several limitations when applied to user review data, which can be broadly categorized into four aspects. First, most studies on Aspect-based Analysis focus on specific domains such as electronics or restaurant reviews [9], making it difficult to ensure that the proposed methods show the same performance and validity in real service environments where a wider range of domains must be addressed. In other words, studies conducted within limited domains do not sufficiently demonstrate multi-domain applicability, which restricts their practical use across industries. Second, Aspect-based Analysis research tends to focus on quantitative metrics such as accuracy and F1-score [35], and lacks qualitative research verifying whether the aspects extracted by the model are semantically valid and practically effective in the real world. As a result, it becomes difficult to determine whether the extracted aspects are trustworthy enough to be used for problem diagnosis or product improvement in practice, which again acts as a limitation in applying the methods in industry. Third, most Aspect-based Analysis research adopts approaches that fine-tune labeled data or assume specific data formats, which leads to high labeling costs and necessitates format conversion [34]. When relying solely on labeled data, it becomes difficult to accurately capture aspects from newly appearing review texts that are not labeled, leading to limitations in responding in real time in actual service environments. Lastly, another issue is that most Aspect-based Analysis studies are conducted mainly on English data. Since large-scale review data is accumulated in various languages in global service environments, models trained only in a specific language face limitations in expanding to multilingual data. As a result, reviews written in newly emerging or underrepresented languages may not be sufficiently analyzed, leading to missed opportunities for satisfaction analysis and service improvement in various markets.
@@ 2.1 Aspect-based Analysis
Research focusing on the analysis of specific aspects in various types of text data has been actively conducted. This technique is utilized in diverse fields such as consumer review analysis, public opinion monitoring, stock market trend prediction [2, 1, 3], and online marketing strategy development [5, 4].
@@
Early studies mainly relied on manually defined rule-based approaches, applying topic modeling methods such as Latent Dirichlet Allocation (LDA) to extract aspect terms from text [6]. [7] proposed extracting aspects using part-of-speech tagging and word frequency data. Another method involved setting preprocessing rules, performing frequency analysis on text data, and manually assigning keywords to predefined categories based on domain knowledge to evaluate keyword importance per attribute [8]. [9] proposed a method to extract aspect terms and opinion terms simultaneously, but since it mainly focused on identifying individual elements, it had limitations in effectively extracting aspect-opinion pairs. These approaches depend on predefined rules, making them difficult to apply to various domains, and because aspect categories are selected based on domain knowledge, they have limitations in reflecting domain-specific terms. Moreover, since most analyses are limited to nouns, it is difficult to conduct sentiment analysis that reflects context.
@@
To overcome these limitations, supervised learning–based approaches have been introduced. Early studies explored the use of Conditional Random Fields (CRFs) to simultaneously extract structural features of sentences and polarity opinions [10]. More recently, deep neural network–based approaches have shown promising performance. [11] applied an attention-based Long Short-Term Memory (LSTM) model for aspect-level sentiment classification, allowing the model to focus on important words in the sentence based on the input aspect. [12] defined a method called Target-oriented Opinion Words Extraction (TOWE) using Inward-Outward LSTM to extract opinion words related to specific targets. Additionally, the RNCRF model, which combines Recurrent Neural Networks (RNNs) and CRFs, was introduced to learn the relationship between aspects and opinions, and a joint learning method was proposed by optimizing CRF parameters [13]. Topic-level sentiment analysis research has also been conducted. Approaches using Online Latent Semantic Indexing and LSTM-based models have been proposed to classify sentiment for specific topics in sentences [14]. [15] conducted research on learning word representations using neural network–based language models. However, supervised deep learning models require a large amount of labeled data, which involves high cost and time for data construction. Moreover, models trained in a specific domain are difficult to apply directly to other applications.
@@
To address these limitations, recent studies have drawn attention to unsupervised learning–based approaches, which aim to perform effective aspect extraction and sentiment analysis without labeled data. Local topic models at the sentence level have been proposed to automatically infer aspects and generate seed sets of sentiment words without manual annotation [16]. [17] proposed a semi-supervised topic model that jointly models aspects and sentiment. [18] combined topic modeling and maximum entropy (MaxEnt) models to separate aspect and opinion words using a semi-supervised approach. More recently, unsupervised aspect term extraction using attention-based Deep Neural Networks (DNNs) has been proposed [19]. These approaches have the advantage of automatically clustering semantically similar aspects without labeled data. However, existing unsupervised methods lack sufficient verification of cross-domain generalization performance, and due to instability during training, they often suffer from low reproducibility and inconsistent performance.
@@
Most previous studies have been conducted in single-language settings. For example, in Aspect-Based Sentiment Analysis using Korean data, [20] explored Attribute Category Sentiment Classification (ACSC) using Bidirectional Encoder Representations from Transformers (BERT). They replaced the [CLS] token with output vectors from tokens related to attribute categories and analyzed performance differences according to input configurations (QA/NLI). In a study using beauty commerce reviews, Importance-Performance Analysis (IPA) and DistilBERT-based sentiment analysis were used to identify areas for product improvement [8]. Another study analyzed sentiment for specific evaluation categories such as food, price, service, and atmosphere using Korean restaurant reviews and proposed a model for predicting detailed ratings [21]. [22] applied attribute-based sentiment analysis using BERT in a movie recommendation system to classify reviews by attributes (e.g., direction, actors, story) and develop a Movie Multi-Criteria recommendation model reflecting user preferences. These studies demonstrate that Aspect-Based Sentiment Analysis enables more fine-grained evaluation of products and services and can enhance the performance of recommendation systems. However, they still lack sufficient verification of generalization across domains and studies considering multilingual applicability remain scarce.
@@ 2.2 Attention Mechanism in Aspect-based Analysis
Recently, attention mechanisms have been widely used in Natural Language Processing (NLP), significantly enhancing the performance of deep neural network models. In particular, pre-trained models such as BERT, Robustly Optimized BERT Pretraining Approach (RoBERTa) effectively capture contextual information through attention, achieving state-of-the-art performance in various NLP tasks. Attention mechanisms learn by focusing on important information within the text. For example, in sequence tagging tasks, they highlight specific words within context [23], and in sequence-to-sequence learning, they help maintain core information during sentence transformation [24]. In sentiment analysis, attention helps models perform more precise predictions by focusing on words related to specific aspects [25], and in machine translation, it contributes to improving semantic alignment between source and target languages [37]. Many researchers have explored the use of attention mechanisms in supervised models for effective Aspect-Based Sentiment Analysis (ABSA) [38]. Additionally, unsupervised attention-based models have been proposed to extract aspects from text without large-scale labeled data [19]. However, most of these existing studies focus on single languages and specific domains, and there is still a lack of research verifying generalization performance in multilingual and multi-domain environments.
@@ 3 Dataset
For Korean review data, we used the Korean String Processing Suite (KSS) [^1] https://github.com/hyunwoongko/kss to segment the text into individual sentences for sentence-level analysis. As a result, we obtained 313,333 sentence-level entries for hotel reviews and 288,874 for food reviews. After removing unnecessary special characters and extra spaces, we refined the text using the Open Korean Text (OKT) morphological analyzer from Konlpy [26]. Since Korean features various grammatical transformations including verb and adjective inflections, we lemmatized these parts of speech and removed case particles (e.g., Korean case particles such as i/ga (subject), eul/reul (object), e (to/at), eseo (from/in)) to ensure analytical consistency. For example, the Korean expressions “맛있었다” (delicious + past tense) and “맛있을 것이다” (delicious + future tense) were both normalized to “맛있다” (delicious). This preprocessing step was critical for reducing variability caused by Korean-specific tense and ending changes that could interfere with model training. We also customized the OKT user dictionary by adding neologisms, slang, and domain-specific proper nouns. This helped prevent incorrect token splits of words not included in the default dictionary and improved the performance of the analyzer. Both preprocessed and original sentences were retained to enable flexible follow-up analyses without data loss.
@@ 4.1.1 Embedding Vector Extraction
We train word embeddings reflecting semantic similarity using the Word2Vec model [15]. Specifically, we apply the Continuous Bag of Words (CBOW) approach, which predicts a center word based on its surrounding context words. For example, in the sentence “The hotel has great service and staff”, if “service” is the center word and the window size is 1, the context words are “great” and “and”.
@@ 4.1.2 Unsupervised Generation of Aspect Category Candidate
Each cluster centroid $c_{k}$ obtained by the K-means algorithm is computed as the mean of the word embeddings $e_{w_{i}}$ belonging to cluster $C_{k}$, and is subsequently $L_{2}$-normalized into a unit vector $\hat{c_{k}}$, as shown in Eq. 2. In this study, the set of normalized centroids $^{\hat{c_{1}},\hat{c_{2}},\dots,\hat{c_{K}}}$ is concatenated into a single matrix, which serves as the initial value of the aspect embedding matrix $E_{c}$. This design enables the semantic structure captured during clustering to be directly transferred to the aspect representation learning stage
@@ 4.2.1 Multi-Head Attention-Based Sentence Representation
Attention mechanisms learn the relationships between words in a sentence and reflect the importance of each word in context. Unlike earlier studies [19] that relied only on attention, we use Multi-Head Attention to learn sentence representations that capture contextual information.
@@
$W^{O}$ is the output weight matrix that projects the concatenated outputs of all attention heads into the final output dimension. Here, headhead denotes an individual attention head, and each attention head is computed as shown in Eq. 4.
@@ 4.2.2 Aspect-Aware Embedding Vector
In this section, the aspect embedding matrix $E_{c}$ used to compute the aspect-aware embedding vector is initialized with the $L_{2}$-normalized centroids obtained from K-means clustering, as described in Section 4.1.2. During training, $E_{c}$ is further refined through the optimization process described in Section 4.2.3, allowing the initial unsupervised clusters to adapt to the learned sentence representations. In this study, the aspect-aware embedding vector $r_{i}$ is computed as shown in Eq. 5, by leveraging the aspect embedding matrix $E_{c}$ calculated in Section 4.1 and the probability vector $p_{t}$, which represents the likelihood of the sentence being associated with each aspect category.
@@ 4.2.3 Optimization of Aspect-Aware Representations
As described in Section 4.2.2, the aspect-aware embedding vector $r_{i}$ is computed using the aspect embedding matrix $E_{c}$ and the probability vector $p_{t}$. In this section, we describe how $E_{c}$ and its related parameters are optimized during training. Specifically, we adopt the Max-Margin Loss to train the aspect-aware embedding representations. In neural language models, the negative sampling technique is widely used to encourage the model to increase the similarity between a sentence and its corresponding aspect category, while reducing the similarity with negative samples. During the application of the negative sampling method, the model is guided to learn more generalizable patterns by minimizing the inner product between the sentence vector $z_{i}$, generated in Section 4.2.1, and negative samples, while maximizing the similarity with the most relevant aspect. The negative samples $z_{n}$ are composed of representation vectors of sentences unrelated to the input sentence, which helps prevent overfitting to a specific aspect during training. To achieve this, we define the hinge loss function as shown in Eq. 6.
@@
Here, $\langle~,$$\rangle$ denotes the inner product between two vectors, and $m$ is a hyperparameter representing the number of negative samples selected for each sentence. To enhance the generalization performance of the model, all vectors are $L_{2}$-normalized to eliminate the influence of vector magnitude. By training the model based on this loss function, it increases the likelihood of assigning a sentence to the correct aspect while reducing similarity to negative samples. This process minimizes the correlation between different aspects, enabling the model to learn more distinguishable aspect representations. Ultimately, the model is trained to minimize the Max-Margin Loss, which guides each input sentence to be assigned to the most appropriate aspect category. Through negative sampling, the model can effectively extract aspect categories and aspect terms across diverse sentence structures, enabling more fine-grained, context-aware analysis.
@@ 5 Experimental Result
This section comprehensively evaluates the performance and practicality of the proposed MUSCAD framework through a series of experiments. First, we empirically determined the optimal number of clusters for each domain during the construction of the automatically labeled dataset using the framework. The number of aspect categories per domain was also determined through expert review. Then, we conducted multi-label classification experiments using the constructed datasets to compare the performance of fine-tuned classification models with that of LLM–based labeling. In addition, a Human Evaluation was conducted to assess whether the automatically generated labels are semantically valid in real-world domain settings. To further verify the effectiveness of MUSCAD’s aspect representation learning module, we conducted comparative experiments with representative unsupervised aspect extraction methods. The comparison was performed using topic coherence (NPMI, UMass), diversity, and embedding coherence metrics to comprehensively evaluate the quality of unsupervised aspect extraction results. Through these experiments, we verified the reliability, scalability, and overall quality of the proposed framework.
@@ 5.1 Experimental Settings
To validate the quality of the labeled datasets generated by the framework, we fine-tuned several pre-trained language models and evaluated their classification performance. We randomly undersampled 40,000 samples for each domain. For the Korean datasets (Hotel and Food), we used KoELECTRA [27], KR-BERT [28], KLUE-BERT [29], Multilingual BERT [30], and XLM-RoBERTa [36]. For the English dataset (Beauty), we used Multilingual BERT, RoBERTa [31], BERT-base, BERT-large [32], XLM-RoBERTa, and DeBERTa-base [33]. All pre-trained language models were fine-tuned using the same hyperparameter settings: a batch size of 16, a maximum token length of 128, and a learning rate of 5e-5. We used the AdamW optimizer and trained each model for 5 epochs. During training, we applied 5-fold cross-validation to ensure stable performance across the dataset and evaluated the models using both Micro-F1 and Macro-F1 scores.